# Thiruvananthapuram 695551, India

### v2

@@ Abstract
We investigate the fully supersymmetric solutions in $\mathcal{N}=3$ and $\mathcal{N}=4$ higher derivative Poincaré supergravity theories. We work in the superconformal formalism and show that flat space is the only fully supersymmetric solution in these theories. However, in $\mathcal{N}=2$ Poincaré supergravity, three fully supersymmetric solutions exist: Bertotti-Robinson geometry ($AdS_{2}\times S^{2}$), flat space and pp wave spacetimes, which are known in the literature. We discuss the reason behind the richer vacua in $\mathcal{N}=2$ supergravity compared to its $\mathcal{N}=3$ and $\mathcal{N}=4$ cousins.
@@ 1 Introduction
[…]
@@
The first classification of supersymmetric solutions using Killing spinor equations was performed by Tod in Tod:1983pm in the context of pure $\mathcal{N}=2$ two-derivative supergravity. The fully supersymmetric solutions are Bertotti-Robinson geometry $(AdS_{2}\times S^{2})$, Minkowski space and pp-wave spacetimes Kowalski-Glikman:1985fjy ; LopesCardoso:1998tkj . In Mohaupt:2000mj ; LopesCardoso:2000qm , the authors considered a class of higher derivative $\mathcal{N}=2$ theories where the higher derivative corrections are encoded in a holomorphic prepotential function and showed that the only $\mathcal{N}=2$ fully supersymmetric stationary solutions are $AdS_{2}\times S^{2}$ and flat space. In this theory, the extremal black holes, which preserves $1/2$ of the supersymmetries, are of the form of Israel-Wilson-Perjés (IWP) metric and it approaches $AdS_{2}\times S^{2}$ geometry at the horizon and Minkowski space at asymptotic infinity LopesCardoso:2000qm . The enhancement of supersymmetry at the horizon led to the discovery of attractor mechanism in the context of $\mathcal{N}=2$ black holes Ferrara:1996dd . For dimension $d>4$, a plethora of solutions appear in supergravity theories with $8$ supercharges, that include black holes, black rings, black branes, Godel type metric, etc. Gauntlett:2002nw ; Gutowski:2003rg ; Chamseddine:2003yy ; Figueroa-OFarrill:2002ecq .
@@
For $\mathcal{N}=4$, $d=4$ supergravity, supersymmetric solutions include the Super-Israel-Wilson-Perjés (SIWP) solutions and pp waves spacetime, but none of these solutions are fully supersymmetric Tod:1995jf ; Bergshoeff:1996gg ; Alonso-Alberca:2002efr ; Bellorin:2005zc . From a string theory point of view, $\mathcal{N}=4$ supergravity can be obtained from compactification of $\mathcal{N}=1,d=10$ heterotic string theory on $T^{6}$ or compactification of $\mathcal{N}=2,d=10$ type-IIA string theory on $K3\times T^{2}$ Chamseddine:1980cp . These parent theories have a flat Minkowski space as the only maximally supersymmetric solution Figueroa-OFarrill:2002ecq . Furthermore, it has recently been argued in Chen:2024gmc that $AdS_{2}\times S^{2}$, which typically appears as a near horizon geometry of extremal supersymmetric black holes, may not possess four Killing spinors (16 supersymmetries). In four dimensions, no solution except flat space has been found to preserve all the 16 supersymmetries. Therefore it remains a conjecture that flat Minkowski space is the only maximally supersymmetric solution in four-dimensional $\mathcal{N}=4$ supergravity.
@@
In this paper, we prove this conjecture for $\mathcal{N}=4$ supergravity in presence of higher derivative corrections. We also show that flat space remains the only maximally supersymmetric solution even for $\mathcal{N}=3$ supergravity in the presence of higher derivative corrections. We also discuss the reason behind the completely different structure of supersymmetric vacua for $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity in contrast with $\mathcal{N}=2$ supergravity which admits richer vacua. We work in the framework of conformal supergravity. Conformal supergravity Fradkin:1985am is a supersymmetric extension of conformal gravity or, equivalently, the conformal extension of Poincaré supergravity. These theories possess a large number of local symmetries, namely diffeomorphism, local Lorentz transformations ($M$), dilatations ($D$), special conformal transformations ($K$), R-symmetries, ordinary supersymmetry (which we will often refer to as $Q$-supersymmetry), and special supersymmetry (often referred to as $S$-supersymmetry). A brief outline of the superconformal formalism is as follows Freedman:2012zz :
@@
  • To close the algebra off-shell, one needs to introduce extra covariant fields, which together with the independent gauge fields form the so-called Weyl multiplet. The superconformal algebra is modified to a soft superconformal algebra where the structure constants become structure functions which depend on the extra covariant matter fields. These extra covariant fields are sometimes referred to as “auxiliary fields”.
@@ 20 $3/2$ 1/2-1/2
$\phi_{ij}$ $\phi^{ij}=(\phi_{ij})^{*}=-\tfrac{1}{2}\varepsilon^{ijkl}\phi_{kl}$ 6 1 0
@@
We now provide a very brief outline of the construction of $\mathcal{N}=4$ higher derivative Poincaré supergravity:
@@ 3 $\mathcal{N}=4$ Fully Supersymmetric Solution
In this section, we find the fully supersymmetric solution in $\mathcal{N}=4$ higher derivative Poincaré supergravity. A fully supersymmetric $\mathcal{N}=4$ solution implies that the background geometry possesses four covariantly constant Killing spinors ($16$ conserved supercharges). We assume the existence of such spinors and demand the $Q$- supersymmetry transformation of the fermions to vanish. The vanishing $Q$-supersymmetry variations of the fermions put stringent conditions on the bosonic fields. We have to find field configurations that satisfy these conditions.
@@
This spinor transforms under under $Q$- and $S$- supersymmetry as
@@
Based on our analysis in this section, we can conclude that there is a unique solution in $\mathcal{N}=4$ supergravity with or without higher derivative corrections that is fully supersymmetric. In this solution, the underlying geometry is a flat space with no electromagnetic fluxes. The coset scalars and the $6n_{v}$ vector multiplet scalars take arbitrary constant values in this background. We summarize the result in Table 4.
@@ 4 $\mathcal{N}=3$ Higher Derivative Supergravity
In this section, we review four dimensional $\mathcal{N}=3$ supergravity using the framework of conformal supergravity. For a detailed construction of $\mathcal{N}=3$ conformal supergravity, see vanMuiden:2017qsh ; Hegde:2018mxv ; Hegde:2021rte . Using these results, higher derivative $\mathcal{N}=3$ Poincaré supergravity was constructed in Hegde:2022wnb . The crucial ingredients are the $\mathcal{N}=3$ Weyl and vector multiplets. A Poincaré supergravity coupled to $n_{V}$ vector multiplets is realised as a Weyl multiplet coupled to $n_{V}+3$ vector multiplets in conformal supergravity where the three compensating vector multiplets have a wrong sign kinetic term. As explained in the introduction, one gauge fixes additional symmetries using the compensators and solves the auxiliary field equations to write the action in Poincaré supergravity variables. In Hegde:2022wnb these further steps were carried out only for the case of pure supergravity. However, for our analysis it is not necessary to eliminate the auxiliary fields since our supersymmetry analysis for the solution is best done in off-shell conformal supergravity variables. We will proceed therefore by considering fully supersymmetric solutions corresponding to the action in conformal supergravity with $n_{V}+3$ vector multiplets, which is equivalent to higher derivative matter coupled Poincaré supergravity.
The $\mathcal{N}=3$ Weyl multiplet is an off-shell multiplet and contains superconformal gauge and matter fields that describe $64+64$ bosonic and fermionic off-shell degrees of freedom. The independent gauge fields include the vierbein $e_{\mu}^{a}$, the gravitini $\psi_{\mu}^{i}$ ($Q$-supersymmetry gauge fields), the $SU(3)$ gauge field ${{V_{\mu}}^{i}}_{j}$ and the dilatation gauge field $b_{\mu}$. The composite or dependent gauge fields are the spin connection ${\omega_{\mu}}^{ab}$, the gauge field associated with the special conformal transformation $f_{\mu}^{a}$, the chiral $U(1)$ gauge field $A_{\mu}$ and $S$-supersymmetry gauge field $\phi_{\mu}^{i}$, where $i,j=1,2,3$.
@@
Apart from the gauge fields, the Weyl multiplet contains several covariant matter fields: both bosonic as well as fermionic. Table 5 summarizes all the fields of the Weyl multiplet along with their algebraic restrictions, $SU(3)$ representations, Weyl weight $w$ under local dilatations and $U(1)$ chiral weight $c$. We follow the notations and conventions of Hegde:2022wnb .
@@ 5 $\mathcal{N}=3$ Fully Supersymmetric Solution
In this section, we find the fully supersymmetric solution in $\mathcal{N}=3$ higher derivative Poincaré supergravity. We will carry this out similar to the analysis of $\mathcal{N}=4$ supersymmetric solutions that we carried out in section-3. We will consider $S$-invariant combintations of fermions and set their variations to zero.
@@
Thus, we obtain the flat space again as the unique fully supersymmetric solution in higher derivative matter coupled $\mathcal{N}=3$ Poincaré supergravity. The scalars can take arbitrary constant values, while the metric remains flat and the fluxes are zero. The result is summarized below in Table-7.
@@ 6 Reason for a richer supersymmetric vacua in $\mathcal{N}=2$ supergravity
$\mathcal{N}=2$ Poincaré supergravity coupled to $n_{V}$ vector multiplets is gauge equivalent to $\mathcal{N}=2$ conformal supergravity coupled to $(n_{V}+1)$ multiplets. As discussed in the introduction, three fully supersymmetric solutions exist in $\mathcal{N}=2$ Poincaré supergravity. The non-stationary fully supersymmetric solutions are pp wave spacetimes. The fully supersymmetric stationary solutions include the $AdS_{2}\times S^{2}$ geometry and the Minkowski flat space. Even in the presence of higher derivative corrections, these two solutions remain the only fully supersymmetric stationary solutions in $\mathcal{N}=2$ Poincaré supergravity. In contrast, $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity have flat space as the unique fully supersymmetric solution.
@@ 7 Conclusions and future directions
In this paper, we used the framework of conformal supergravity to show that flat space is the only fully supersymmetric solution in the matter coupled $\mathcal{N}=4$ and $\mathcal{N}=3$ higher derivative Poincaré supergravity. We have also argued that the vanishing of the auxiliary fields $T_{ab}^{ij}$ and $T_{ab}^{i}$ is the sole reason for getting flat space as a unique fully supersymmetric solution in $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity respectively. Our results hold to all order in the derivative expansion for a class of supergravity Lagrangians that are constructed using the conformal supergravity Lagrangian of Butter:2019edc ; Hegde:2022wnb followed by gauge fixing the additional super-conformal symmetries. However, the input provided by the details of the conformal supergravity action for our results is very limited. In particular, only the $D$-field equation of motion has been used from the conformal supergravity action to arrive at our results. Therefore we expect our results to hold more generally for $\mathcal{N}=3,4$ Poincaré supergravity to all order in the derivative expansion.
@@ Acknowledgements.
We thank Ashoke Sen and Amitabh Virmani for discussions. AB and BS thank Harish Chandra Research Institute, Prayagraj for its hospitality during the initial stages of this project. AB thanks organizers of National Strings Meeting 2024, IIT Ropar where this work was presented. SH is funded by the European Union (ERC, UNIVERSE PLUS, 101118787). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v3

@@ Abstract
We investigate the fully supersymmetric solutions in a class of $\mathcal{N}=3$ and $\mathcal{N}=4$ higher derivative Poincaré supergravity theories. These class of theories are obtained within the framework of conformal supergravity using the standard Weyl multiplet and contains a set of terms related to the Weyl square term by supersymmetry. We work in the superconformal formalism and show that flat spacetime is the only fully supersymmetric solution in these theories. However, in $\mathcal{N}=2$ Poincaré supergravity theories that falls into the same class, two fully supersymmetric stationary solutions exist: Bertotti-Robinson geometry ($AdS_{2}\times S^{2}$) and flat spacetime, which are known in the literature. We discuss the reason behind the richer vacua in $\mathcal{N}=2$ supergravity compared to its $\mathcal{N}=3$ and $\mathcal{N}=4$ cousins.
@@ 1 Introduction
The spinors in the subset are known as Killing spinors and the schematic equations in (1.1) are the Killing Spinor equations. For classical field configurations, the fermions vanish, and the first condition of (1.1) is trivially satisfied. A fully supersymmetric solution supports the maximal number of Killing spinors, i.e., $M=\mathcal{N}$ ( $4\mathcal{N}$ conserved supercharges). [^2] In four spacetime dimensions, the basic spinors are Majorana spinors with four real components. For spinors in other dimensions, look at Freedman:2012zz . For $M<\mathcal{N}$, the solution is said to preserve $M/\mathcal{N}$ of total supersymmetries. The Killing spinor equations serve a dual purpose. Firstly, for a given solution, one can find the Killing spinors by solving the vanishing gravitini field variations
@@
The first classification of supersymmetric solutions using Killing spinor equations was performed by Tod in Tod:1983pm in the context of pure $\mathcal{N}=2$ two-derivative supergravity. The fully supersymmetric solutions are Bertotti-Robinson geometry $(AdS_{2}\times S^{2})$, Minkowski space and pp-wave spacetimes Kowalski-Glikman:1985fjy ; LopesCardoso:1998tkj . In Mohaupt:2000mj ; LopesCardoso:2000qm , the authors considered a class of higher derivative $\mathcal{N}=2$ theories where the higher derivative corrections are encoded in a holomorphic prepotential function and showed that the only $\mathcal{N}=2$ fully supersymmetric stationary solutions are $AdS_{2}\times S^{2}$ and flat spacetime. In this theory, the extremal black holes, which preserves $1/2$ of the supersymmetries, are of the form of Israel-Wilson-Perjés (IWP) metric and it approaches $AdS_{2}\times S^{2}$ geometry at the horizon and Minkowski space at asymptotic infinity LopesCardoso:2000qm . The enhancement of supersymmetry at the horizon led to the discovery of attractor mechanism in the context of $\mathcal{N}=2$ black holes Ferrara:1996dd . For dimension $d>4$, a plethora of solutions appear in supergravity theories with $8$ supercharges, that include black holes, black rings, black branes, Godel type metric, etc. Gauntlett:2002nw ; Gutowski:2003rg ; Chamseddine:2003yy ; Figueroa-OFarrill:2002ecq .
@@
For $\mathcal{N}=4$, $d=4$ supergravity, supersymmetric solutions include the Super-Israel-Wilson-Perjés (SIWP) solutions and pp waves spacetime, but none of these solutions are fully supersymmetric Tod:1995jf ; Bergshoeff:1996gg ; Alonso-Alberca:2002efr ; Bellorin:2005zc . From a string theory point of view, $\mathcal{N}=4$ supergravity can be obtained from compactification of $\mathcal{N}=1,d=10$ heterotic string theory on $T^{6}$ or compactification of $\mathcal{N}=2,d=10$ type-IIA string theory on $K3\times T^{2}$ Chamseddine:1980cp . These parent theories have a flat Minkowski space as the only maximally supersymmetric solution Figueroa-OFarrill:2002ecq . Furthermore, it has recently been argued in Chen:2024gmc that $AdS_{2}\times S^{2}$, which typically appears as a near horizon geometry of extremal supersymmetric black holes, may not possess four Killing spinors (16 supersymmetries). In four dimensions, no solution except flat spacetime has been found to preserve all the 16 supersymmetries. Therefore it remains a conjecture that flat Minkowski space is the only maximally supersymmetric solution in four-dimensional $\mathcal{N}=4$ supergravity.
@@
In this paper, we prove this conjecture for $\mathcal{N}=4$ supergravity in the presence of a class of higher derivative corrections that is obtained within the framework of conformal supergravity using the “standard Weyl multiplet” (See footnote-3.) and contains all terms related to the Weyl squared terms by supersymmetry. We also show that flat spacetime remains the only maximally supersymmetric solution even for $\mathcal{N}=3$ supergravity in the presence of the same class of higher derivative corrections. We also discuss the reason behind the completely different structure of supersymmetric vacua for $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity in contrast with $\mathcal{N}=2$ supergravity which admits richer vacua. We work in the framework of conformal supergravity. Conformal supergravity Fradkin:1985am is a supersymmetric extension of conformal gravity or, equivalently, the conformal extension of Poincaré supergravity. These theories possess a large number of local symmetries, namely diffeomorphism, local Lorentz transformations ($M$), dilatations ($D$), special conformal transformations ($K$), R-symmetries, ordinary supersymmetry (which we will often refer to as $Q$-supersymmetry), and special supersymmetry (often referred to as $S$-supersymmetry). A brief outline of the superconformal formalism is as follows Freedman:2012zz :
@@
  • To close the algebra off-shell, one needs to introduce extra covariant fields, which together with the independent gauge fields form the so-called Weyl multiplet. The superconformal algebra is modified to a soft superconformal algebra where the structure constants become structure functions which depend on the extra covariant matter fields. These extra covariant fields are sometimes referred to as “auxiliary fields”.  [^3] There are different ways of introducing the auxiliary fields which leads to different variants of Weyl multiplets: Standard and Dilaton. In this paper we confine ourselves to the standard Weyl multiplet.
@@ 20 $3/2$ 1/2-1/2
$\phi_{ij}$ $\phi^{ij}\equiv(\phi_{ij})^{*}=-\tfrac{1}{2}\varepsilon^{ijkl}\phi_{kl}$ 6 1 0
@@
We now provide a very brief outline of the construction of $\mathcal{N}=4$ higher derivative Poincaré supergravity within the framework of conformal supergravity constructed using the standard Weyl multiplet. Such theories contains a class of higher derivative terms that contains all terms related to Weyl square terms by supersymmetry:
@@ 3 $\mathcal{N}=4$ Fully Supersymmetric Solution
In this section, we find the fully supersymmetric solution in $\mathcal{N}=4$ higher derivative Poincaré supergravity that is considered in the previous section. A fully supersymmetric $\mathcal{N}=4$ solution implies that the background geometry possesses four covariantly constant Killing spinors ($16$ conserved supercharges). We assume the existence of such spinors and demand the $Q$- supersymmetry transformation of the fermions to vanish. The vanishing $Q$-supersymmetry variations of the fermions put stringent conditions on the bosonic fields. We have to find field configurations that satisfy these conditions.
@@
This spinor transforms under $Q$- and $S$- supersymmetry as
@@
Based on our analysis in this section, we can conclude that there is a unique solution in $\mathcal{N}=4$ supergravity with or without higher derivative corrections that is fully supersymmetric. In this solution, the underlying geometry is a flat spacetime with no electromagnetic fluxes. The coset scalars and the $6n_{v}$ vector multiplet scalars take arbitrary constant values in this background. We summarize the result in Table 4.
@@ 4 $\mathcal{N}=3$ Higher Derivative Supergravity
In this section, we review four dimensional $\mathcal{N}=3$ supergravity using the framework of conformal supergravity. For a detailed construction of $\mathcal{N}=3$ conformal supergravity, see vanMuiden:2017qsh ; Hegde:2018mxv ; Hegde:2021rte . Using these results, a class of higher derivative $\mathcal{N}=3$ Poincaré supergravity was constructed in Hegde:2022wnb , that contains all the terms related to Weyl squared term by supersymmetry. The crucial ingredients are the $\mathcal{N}=3$ Weyl and vector multiplets. A Poincaré supergravity coupled to $n_{V}$ vector multiplets is realised as a Weyl multiplet coupled to $n_{V}+3$ vector multiplets in conformal supergravity where the three compensating vector multiplets have a wrong sign kinetic term. As explained in the introduction, one gauge fixes additional symmetries using the compensators and solves the auxiliary field equations to write the action in Poincaré supergravity variables. In Hegde:2022wnb these further steps were carried out only for the case of pure supergravity. However, for our analysis it is not necessary to eliminate the auxiliary fields since our supersymmetry analysis for the solution is best done in off-shell conformal supergravity variables. We will proceed therefore by considering fully supersymmetric solutions corresponding to the action in conformal supergravity with $n_{V}+3$ vector multiplets, which is equivalent to higher derivative matter coupled Poincaré supergravity.
The $\mathcal{N}=3$ Weyl multiplet is an off-shell multiplet and contains superconformal gauge and matter fields that describe $64+64$ bosonic and fermionic off-shell degrees of freedom. The independent gauge fields include the vierbein $e_{\mu}^{a}$, the gravitini $\psi_{\mu}^{i}$ ($Q$-supersymmetry gauge fields), the $SU(3)$ gauge field ${{V_{\mu}}^{i}}_{j}$ and the dilatation gauge field $b_{\mu}$. The composite or dependent gauge fields are the spin connection ${\omega_{\mu}}^{ab}$, the gauge field associated with the special conformal transformation $f_{\mu}^{a}$, the chiral $U(1)$ gauge field $A_{\mu}$ and $S$-supersymmetry gauge field $\phi_{\mu}^{i}$, where $i,j=1,2,3$. Apart from the gauge fields, the Weyl multiplet contains several covariant matter fields: both bosonic as well as fermionic. Table 5 summarizes all the fields of the Weyl multiplet along with their algebraic restrictions, $SU(3)$ representations, Weyl weight $w$ under local dilatations and $U(1)$ chiral weight $c$. We follow the notations and conventions of Hegde:2022wnb .
@@ 5 $\mathcal{N}=3$ Fully Supersymmetric Solution
In this section, we find the fully supersymmetric solution in $\mathcal{N}=3$ higher derivative Poincaré supergravity considered in the previous section. We will carry this out similar to the analysis of $\mathcal{N}=4$ supersymmetric solutions that we carried out in section-3. We will consider $S$-invariant combinations of fermions and set their variations to zero.
@@
Thus, we obtain the flat spacetime again as the unique fully supersymmetric solution in higher derivative matter coupled $\mathcal{N}=3$ Poincaré supergravity. The scalars can take arbitrary constant values, while the metric remains flat and the fluxes are zero. The result is summarized below in Table-7.
@@ 6 Reason for a richer supersymmetric vacua in $\mathcal{N}=2$ supergravity
$\mathcal{N}=2$ Poincaré supergravity coupled to $n_{V}$ vector multiplets is gauge equivalent to $\mathcal{N}=2$ conformal supergravity coupled to $(n_{V}+1)$ multiplets. As discussed in the introduction, three fully supersymmetric solutions exist in $\mathcal{N}=2$ Poincaré supergravity. The non-stationary fully supersymmetric solutions are pp wave spacetimes. The fully supersymmetric stationary solutions include the $AdS_{2}\times S^{2}$ geometry and flat spacetime. Even in the presence of higher derivative corrections, these two solutions remain the only fully supersymmetric stationary solutions in $\mathcal{N}=2$ Poincaré supergravity. In contrast, $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity have flat spacetime as the unique fully supersymmetric solution.
@@ 7 Conclusions and future directions
In this paper, we used the framework of conformal supergravity to show that flat spacetime is the only fully supersymmetric solution in the matter coupled $\mathcal{N}=4$ and $\mathcal{N}=3$ higher derivative Poincaré supergravity that contains a class of higher derivative terms related to the Weyl squared term by supersymmetry. We have also argued that the vanishing of the auxiliary fields $T_{ab}^{ij}$ and $T_{ab}^{i}$ is the sole reason for getting flat spacetime as a unique fully supersymmetric solution in $\mathcal{N}=4$ and $\mathcal{N}=3$ supergravity respectively. Our results hold to all order in the derivative expansion for a class of supergravity Lagrangians that are constructed using the conformal supergravity Lagrangian of Butter:2019edc ; Hegde:2022wnb followed by gauge fixing the additional super-conformal symmetries. However, the input provided by the details of the conformal supergravity action for our results is very limited. In particular, only the $D$-field equation of motion has been used from the conformal supergravity action to arrive at our results. Therefore we expect our results to hold more generally for $\mathcal{N}=3,4$ Poincaré supergravity to all order in the derivative expansion. The class of higher derivative theories considered in this paper are the most general ones within the framework of $\mathcal{N}=4$ and $\mathcal{N}=3$ conformal supergravity theories constructed using the standard Weyl multiplet. However, one may be able to construct more general class of higher derivative theories within the framework of $\mathcal{N}=4$ and $\mathcal{N}=3$ conformal supergravity but instead using the dilaton Weyl multiplets Ciceri:2024xxf ; Adhikari:2024esl ; Adhikari:2024qxg ; Adhikari:2025wwb . It would be interesting to extend our analysis to such class of theories.
@@ Acknowledgements.
We thank Ashoke Sen and Amitabh Virmani for discussions. AB and BS thank Harish Chandra Research Institute, Prayagraj for its hospitality during the initial stages of this project. AB thanks organizers of National Strings Meeting 2024, IIT Ropar where this work was presented. SH is funded by the European Union (ERC, UNIVERSE PLUS, 101118787). BS thanks the organisers of the workshop “New Frontiers in Supergravity, Strings and Related Topics” held at Hefei, China, where this work was presented. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.